OpenStudy (anonymous):

Integral from 0 to a 2e^x(e^x+1)^3 dx = 120. Find X.

OpenStudy (anonymous):

a = ln 3?

OpenStudy (anonymous):

Use u=e^x+1 du=e^x dx That makes your integral int 2u^3 so you have (1/2)(e^x+1)^4 evaluated from 0 to a. Evaluate it, solve for a.